giannamarie:

Elliott’s backyard is 4800 square feet. His dad is going to pave 15% of this area to use as a patio. How many square feet will be paved?

justasoftie:

Welcome to Qc! What about this problem do you not understand?

iosangel:

4800*15% =\[\frac{ 4800*15 }{ 100 }\]= \[\frac{ 72000 }{ 100}\]=720

justasoftie:

Although, you should not post direct and answers because that is not helping, that answer would be correct

justasoftie:

I was going to try to walk you through it, but there's your answer